# WAVE: Learning Unified & Versatile Audio-Visual Embeddings with Multimodal LLM

### v1

@@ Abstract
While embeddings from multimodal large language models (LLMs) excel as general-purpose representations, their application to dynamic modalities like audio and video remains underexplored. We introduce WAVE (unified & versatile audio-visual embeddings), the first LLM-based embedding that creates a unified representation space for text, audio, and video modalities. WAVE employs a novel hierarchical feature fusion strategy and a joint multi-modal, multi-task training approach to enable two key capabilities: any-to-any cross-modal retrieval and the generation of prompt-aware embeddings tailored to user instructions. Experimentally, WAVE sets a new state-of-the-art on the MMEB-v2 video benchmark and achieves superior results in audio and video-to-audio retrieval. Its prompt-aware nature also yields remarkable performance in multimodal question answering, significantly outperforming existing embedding models. Ablation studies validate our joint training strategy, demonstrating improved performance across all modalities. With a newly introduced benchmark for versatile audio-visual learning, WAVE opens up broad possibilities for cross-modal, any-to-any applications. Our code, checkpoints, and data will be released.
@@ 3.1 Model Architecture
After TMRoPE is applied, the multimodal token sequence is fed into the LLM. Inspired by the use of the hidden state of the final “EOS” token by last-token pooling, we aggregate the last output tokens from multiple LLM layers to construct embeddings for non-text modalities. This design captures both low-level perceptual cues and high-level semantic abstractions. A lightweight fusion module—implemented as a two-layer multi-layer perceptron (MLP) with GELU activation (Hendrycks & Gimpel, 2016) is then used to refine and compress the embeddings. For the text-only scenario, we retain the standard last-token pooling approach, which prior studies have shown to be highly effective.
@@ 4.1 Model Specifications
As for the model input, videos, in general, are sampled at 2 frames per second, with a maximum of 128 frames sampled. For videos longer than 64 seconds, 128 frames are uniformly sampled to conserve memory resources. For audio input, the waveform signals are resampled to 16,000 Hz.
@@ 4.2 Training Specifications
MSVD (Chen & Dolan, 2011) (visual, text) 24 K
@@ 4.3 Evaluation Specifications
Details of the evaluation data and metrics are shown in Table 2. All evaluation tasks are formulated as “query-to-target” retrieval. For video-centric tasks, we adopt the video subset from MMEB-v2 (Meng et al., 2025) as our foundation, and we also augment the evaluations on the recent benchmark LoVR (Cai et al., 2025). In the audio domain, our evaluation encompasses both retrieval and QA tasks as well. Beyond these text-centric scenarios, more challenging tasks such as video-to-audio retrieval and video-to-music retrieval are evaluated, which test the model’s ability to map visual semantics to auditory concepts directly within its unified embedding space.
@@ 5.2 Analysis of prompt-aware embeddings
Beyond retrieval, WAVE leverages its LLM backbone’s reasoning to produce prompt-aware embeddings conditioned on textual instructions (see Appendix B for a case study). The ability to follow instructions is crucial for embedding MLLMs in QA tasks, which can be shown by the following question example from Video-MME (Fu et al., 2025): Which of the following features/items is not discussed in the video in relation to the tomb? A. Inkstone. B. Niche. C. Jade. D. Sacrificial table.
@@ Appendix A The Use of Large Language Models
Appendix B Case study of prompt-aware embeddings
@@
Appendix C Analysis of Dual Speech & Audio Encoders

### v2

@@ Abstract
While embeddings from multimodal large language models (LLMs) excel as general-purpose representations, their application to dynamic modalities like audio and video remains underexplored. We introduce WAVE (unified & versatile audio-visual embeddings), the first LLM-based embedding that creates a unified representation space for text, audio, and video modalities. WAVE employs a novel hierarchical feature fusion strategy and a joint multi-modal, multi-task training approach to enable two key capabilities: any-to-any cross-modal retrieval and the generation of prompt-aware embeddings tailored to user instructions. Experimentally, WAVE sets a new state-of-the-art on the MMEB-v2 video benchmark and achieves superior results in audio and video-to-audio retrieval. Its prompt-aware nature also yields remarkable performance in multimodal question answering, significantly outperforming existing embedding models. Ablation studies validate our joint training strategy, demonstrating improved performance across all modalities. With a newly introduced benchmark for versatile audio-visual learning, WAVE opens up broad possibilities for cross-modal, any-to-any applications. Our code and checkpoints are released at https://github.com/TCL606/WAVE.
@@ 3.1 Model Architecture
To structure the multimodal input tokens for the LLM, we employ specific interleaving strategies. For audio-only input, the speech-related and audio event-related tokens, which are equal in number due to identical encoder frequencies, are interleaved on a one-to-one basis to form a unified auditory token sequence. For synchronised audio-visual input, both the visual and auditory token sequences are partitioned into several segments corresponding to the number of sampled frames. These segments are then interleaved to create the audio-visual token sequence. Finally, the text tokens of the prompt are appended to the end to form the input token sequence for the LLM.
@@
After TMRoPE is applied, the multimodal token sequence is fed into the LLM. Inspired by the use of the hidden state of the final “EOS” token by last-token pooling, we aggregate the last output tokens from multiple LLM layers to construct embeddings for non-text modalities. This design captures both low-level perceptual cues and high-level semantic abstractions. A lightweight fusion module—implemented as a two-layer multi-layer perceptron (MLP) with GELU activation (Hendrycks and Gimpel, 2016) is then used to refine and compress the embeddings. For the text-only scenario, we retain the standard last-token pooling approach, which prior studies have shown to be highly effective.
@@ 4.1 Model Specifications
As for the model input, videos, in general, are sampled at 2 frames per second, with a maximum of 128 frames sampled. For videos longer than 64 seconds, 128 frames are uniformly sampled to conserve memory resources. The maximum resolution for each frame is 176400 pixels. For audio input, the waveform signals are resampled to 16,000 Hz. Other preprocessing settings are identical to those of Qwen2.5-Omni.
@@ 4.2 Training Specifications
MSVD (Chen and Dolan, 2011) (visual, text) 24 K
@@ 4.3 Evaluation Specifications
Details of the evaluation data and metrics are shown in Table 2. All evaluation tasks are formulated as “query-to-target” retrieval. For video-centric tasks, we adopt the video subset from MMEB-v2 (Meng et al., 2025) as our foundation, and we also augment the evaluations on the recent benchmark LoVR (Cai et al., 2025). In the audio domain, our evaluation encompasses both retrieval and QA tasks as well. Beyond these text-centric scenarios, more challenging tasks such as video-to-audio retrieval and video-to-music retrieval are evaluated, which test the model’s ability to map visual semantics to auditory concepts directly within its unified embedding space. More details of the inference procedure for each evaluation task are shown in Appendix B.
@@ 5.2 Analysis of prompt-aware embeddings
Beyond retrieval, WAVE leverages its LLM backbone’s reasoning to produce prompt-aware embeddings conditioned on textual instructions (see Appendix D for a case study). The ability to follow instructions is crucial for embedding MLLMs in QA tasks, which can be shown by the following question example from Video-MME (Fu et al., 2025): Which of the following features/items is not discussed in the video in relation to the tomb? A. Inkstone. B. Niche. C. Jade. D. Sacrificial table.
@@ Appendix A The Use of Large Language Models
Appendix B Inference procedure for evaluation tasks
The evaluation tasks can also be divided into two categories: retrieval tasks and QA tasks. Tasks that use a single, unified general prompt to extract embeddings for all test samples are regarded as retrieval tasks. This includes the classification, retrieval, and moment retrieval subsets of MMEB-v2-Video. Conversely, tasks that use separate and specific questions as prompts to extract embeddings are considered QA tasks.
For all retrieval tasks, we generate an embedding for each sample in the test set based on its input video/audio, using the fixed text prompt, ”Please describe the video/audio.” If the task involves retrieving text, the corresponding ground-truth text captions for each sample are also used to generate text embeddings. The entire test set then forms the candidate pool for the retrieval task. The candidate with the highest similarity score to the query embedding is selected as the model’s prediction.
For QA tasks, all QA evaluations for the embedding LLMs are performed using an embedding-based methodology. To be specific, the inference procedure is as follows:
  1. 1.
    The model first generates a single embedding that is conditioned on both the source modality (video/audio) and the provided question text.
  2. 2.
    Separately, the model generates an embedding for the text of each answer option.
  3. 3.
    The similarity between the question-conditioned video/audio embedding and each of the option text embeddings is then calculated.
  4. 4.
    The option with the highest similarity score is selected as the model’s predicted answer.
Appendix C More Evaluation Results of WAVE
In the retrieval task of MMEB-v2-Video, only one direction, text-to-video retrieval, was evaluated. Similarly, for the audio retrieval and audio-visual retrieval tasks presented in Table 4, only a single direction was assessed, i.e, text-to-audio and video-to-audio retrieval, respectively. This focus was chosen because these directions more closely align with practical, real-world application scenarios like multimodal search and recommendation. However, it is undeniable that performance in the reverse direction is also important. In Table 8, we provide supplementary results for video retrieval (V-RET), audio retrieval (A-RET), and audio-visual retrieval (AV-RET) in the other direction. WAVE can also achieve competitive results in the other direction.
Table 8: Results of video-to-text, audio-to-text and audio-to-video retrieval. Corresponding reference models and their scores are also provided.
Method V-RET A-RET AV-RET
MSR-VTT VATEX YouCook2 Clotho VGGSound
Reference Model GME 7B (Zhang et al., 2024) (Mei et al., 2024) encoder-only retrieval model (ours)
Reference Value 32.2 33.0 10.0 27.1 8.2
WAVE 7B 55.1 55.4 29.8 24.5 25.8
Appendix D Case study of prompt-aware embeddings
@@
Appendix E Analysis of Dual Speech & Audio Encoders
@@
Appendix F The Effect of Image Training
WAVE is primarily trained on video data and achieves strong performance on video retrieval-related tasks. However, this success is not due to a narrow specialization in video. In fact, WAVE is not specialized only for video retrieval, and image data actually helps. In a preliminary study, we compared the results of training solely on video data against training on a mixture of video data and a nearly equal amount of image data from the MMEB-v1 training set. This experiment utilized the same pure-visual video retrieval setup as described in Section 5.4. Table 10 presents the results of this comparison.
Table 10: Comparison of results for training with and without image data. We evaluate text-to-image retrieval (Liu et al., 2021) on the VisualNews dataset and text-to-video retrieval on MMEB-v2-Video.
Training Data Image RET MMEB-v2-Video RET
VisualNews Average MSR-VTT VATEX MSVD DiDeMo YouCook2
Video Data 54.9 49.6 52.1 46.2 69.7 53.0 27.2
Video Data + Image Data (MMEB-v1) 74.8 50.2 52.4 46.2 69.0 55.8 27.8
As the results show, including image data substantially improves image retrieval and provides a slight gain in video retrieval. This shows that WAVE’s strong performance is not due to avoiding image tasks or specializing exclusively in video.